User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Arwin

Present: Jamie, Lucia, Tiege, Pat, Jonny, MAS, Liz

Apologies: Lucas, Bryn, Ross

Absent: Jesse

General Issues:

* Jamie is fixing the fringe checking script, should be ready next week sometime. 

Handover Issues:

* Yarragadee elevation drive issues during r4798 (Liz) again. Ongoing issues, best to send it to southpole after doing setup, before starting the schedule. Failing that try your normal resets. Any permanent fix in the future?
- Jamie to talk to Brett and other yg staff.......
* RD1706 (Mas) had issues with MKV being in a weird state and not running DIMino. Warren used ''SU reboot'' to restart it remotely which fixed the problem.
* R1798 (Ross) had an issue where the schedule returned to the start and went through the scans upto present over about 10 mins, missing a few scans. One time issue?
* r4797 (Arwin) Yarragadee has issues with the drives once again, none of the resets fixed it. Brett repeated the same procedure except he waiting a while between turning the drives on and pressing 'operate', with success.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_07_06.txt · Last modified: 2017/07/14 04:38 by Arwin Kahlon